The hour pay of an administrator last year rose from $105/hr to$112/ hr.what is the percent increase
Found 2 solutions by sachi, lynnlo:
Answer by sachi(548)

The hour pay of an administrator last year rose from $105/hr to$112/ hr
the percent increase =(112-105)*100/105=7*100/105=100/15=20/3=6 2/3%

Answer by lynnlo(4176)

6.66% is the increase in pay